# Still More Shades of Null: A Benchmark for Responsible Missing Value Imputation

### v1

@@ Appendix D Additional Experimental Results for Missingness Shift
D.1. Stability of Predictive Model
This section supplements the discussion of Section 6.3. It includes the label stability plots that were omitted in the main body of the paper due to space constraints.
Figure 30. Label Stability of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates performance of the model trained on clean data.
Figure 31. Label Stability of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
D.2. Fixed Train and Test Missingness Fractions
F1 of predictive model. The F1 scores of predictive models trained with various MVI techniques under different train and test missingness conditions, including missingness shift (scenarios S1-9 from Table 1), are presented in Figure 34. The plot demonstrates that the F1 scores of predictive models are sensitive to missingness shifts across different datasets, as also discussed in Section 6. The impact of missingness shift is most significant for MCAR train & MNAR test and MAR train & MNAR test scenarios on the diabetes, german, and heart datasets. This observation aligns with the statement in (Shadbahr et al., 2023) that MNAR is more complex to model than MCAR and MAR. Additionally, miss-forest emerges as the most robust technique in handling missingness shifts across all datasets and settings.
@@ D.2. Fixed Train and Test Missingness Fractions
(d) Deletion
Figure 32. F1 of different models (colors in legend) on the diabetes dataset, for best and worst performing MVI techniques (subplots).
@@
(d) Deletion
Figure 33. Stability of different models (colors in legend) on the diabetes dataset, for best and worst performing MVI techniques (subplots).
@@
Figure 34. F1 of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-9)
@@
Figure 35. Label stability of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-9)
Fairness of predictive model. The fairness of predictive models according to true positive rate difference is illustrated in Figure 36. The plot reveals that the impact of missingness shifts on fairness varies depending on the missingness type of the test set. For instance, for the diabetes dataset, a predictive model trained on an MCAR train set can exhibit worse TPRD on an MNAR test set compared to an MCAR test set. Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set. We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 37-54. These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across all ML and DL-based imputers and all six datasets.
@@
Figure 37. Disparate Impact of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 38. True Negative Rate Difference of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 39. Selection Rate Difference of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 40. Disparate Impact of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 41. True Negative Rate Difference of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 42. Selection Rate Difference of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 43. Disparate Impact of different models (colors in legend) on the folk dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 44. True Negative Rate Difference of different models (colors in legend) on the folk dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 45. Selection Rate Difference of different models (colors in legend) on the folk dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 46. Disparate Impact of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 47. True Negative Rate Difference of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 48. Selection Rate Difference of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 49. Disparate Impact of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 50. True Negative Rate Difference of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 51. Selection Rate Difference of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 52. Disparate Impact of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 53. True Negative Rate Difference of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots).
@@
Figure 54. Selection Rate Difference of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots).
D.3. Variable Train and Test Missingness Fractions
Fairness under train set missingness. Figures 55, 56, and 57 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
@@ D.3. Variable Train and Test Missingness Fractions
Fairness under test set missingness. Figures 58, 59, and 60 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.
@@
The plots reinforce the assertion in Section 6.2 that the trend depends significantly on the test missingness type. For example, the subfigures for MCAR train & MNAR test and MAR train & MNAR test show the largest deviations with increasing test error rates for all three fairness metrics, while plots for other settings remain relatively flat. Interestingly, when null imputers and models are fitted on the MNAR train set, model performance on the MNAR test set across different test error rates do not change significantly compared to MCAR train & MNAR test and MAR train & MNAR test settings. This indicates that the missingness mechanism in the train set can significantly affect model fairness on the MNAR test set. Additionally, miss-forest and auto-ml proved to be the most robust MVI techniques against test set missingness shifts under all conditions.
@@
Figure 55. True Negative Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
@@
Figure 56. Disparate Impact of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
@@
Figure 57. Selection Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
@@
Figure 58. True Negative Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
@@
Figure 59. Disparate Impact of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
@@
Figure 60. Selection Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.

### v2

@@ Appendix D Additional Experimental Results for Missingness Shift
D.1. Fixed Train and Test Missingness Fractions
F1 of predictive model. F1 scores of predictive models trained with various MVI techniques under different train and test missingness conditions, including missingness shift (scenarios S1-9 from Table 1), are presented in Figure 24. The plot demonstrates that F1 scores of predictive models are sensitive to missingness shifts across different datasets, as also discussed in Section 5.1. The impact of missingness shift is most significant for MCAR train & MNAR test and MAR train & MNAR test scenarios on the diabetes, german, and heart datasets. This observation aligns with the statement in (Shadbahr et al., 2023) that MNAR is more complex to model than MCAR and MAR. Additionally, miss-forest emerges as the most robust technique in handling missingness shifts across all datasets and settings.
@@ D.1. Fixed Train and Test Missingness Fractions
(d) MissForest
Figure 22. F1 of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift (Scenarios S1-S9).
@@
(d) MissForest
Figure 23. Stability of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift (Scenarios S1-S9).
@@
(g) Folk Employment (302,640 samples)
Figure 24. F1 of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-S9). The blue line indicates the performance of a model trained on clean data.
@@
(g) Folk Employment (302,640 samples)
Figure 25. Label stability of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-S9). The blue line indicates the performance of a model trained on clean data.
Fairness of predictive model. The fairness of predictive models according to true positive rate difference is illustrated in Figure 26. The plot reveals that model fairness is highly sensitive to missingness shift. For instance, for the diabetes dataset, a predictive model trained on an MCAR train set can exhibit worse TPRD on an MNAR test set compared to an MCAR test set. Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set. We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 27-44. These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across all ML and DL-based imputers and all our datasets.
@@
Figure 27. Disparate Impact of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 28. True Negative Rate Difference of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 29. Selection Rate Difference of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 30. Disparate Impact of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 31. True Negative Rate Difference of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 32. Selection Rate Difference of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 33. Disparate Impact of different models (colors in legend) on the folk-income dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 34. True Negative Rate Difference of different models (colors in legend) on the folk-income dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 35. Selection Rate Difference of different models (colors in legend) on the folk-income dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 36. Disparate Impact of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 37. True Negative Rate Difference of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 38. Selection Rate Difference of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 39. Disparate Impact of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 40. True Negative Rate Difference of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 41. Selection Rate Difference of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 42. Disparate Impact of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 43. True Negative Rate Difference of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
@@
Figure 44. Selection Rate Difference of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
D.2. Variable Train and Test Missingness Rates
Correctness under variable test set missingness. This subsection supplements the discussion of Section 5.1. Figure 45 shows effect of varying test missingness rates on the F1 of the Random Forest model on diabetes, under missingness shifts.
Figure 45. F1 of the Random Forest model on diabetes, as a function of test missingness rate, under different missingness scenarios. The dashed line indicates performance of the model trained on clean data.
Fairness (TPRD) under variable test set missingness. This subsection supplements the discussion of Section 5.2. Figure 46 shows effect of varying test missingness rates on the True Positive Rate Difference (TPRD) of the Random Forest model on diabetes, under missingness shifts.
Stability under variable test set missingness. This subsection supplements the discussion of Section 5.3. It includes the label stability plots that were omitted in the main body of the paper due to space constraints. Figure 47 shows effect of varying test missingness rates on the Label Stability of the Random Forest model on diabetes, under missingness shifts.
Figure 47. Label Stability of the Random Forest model on diabetes, as a function of test missingness rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Other fairness metrics under variable train set missingness. Figures 48, 49, and 50 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
@@ D.2. Variable Train and Test Missingness Rates
Other fairness metrics under variable test set missingness. Figures 51, 52, and 53 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.
@@
The plots reinforce the assertion in Section 5.2 that model fairness is highly sensitive to missingness shift. For example, the subfigures for MCAR train & MNAR test and MAR train & MNAR test show the largest deviations with increasing test error rates for all three fairness metrics, while plots for other settings remain relatively flat. Interestingly, when null imputers and models are fitted on the MNAR train set, model performance on the MNAR test set across different test error rates do not change significantly compared to MCAR train & MNAR test and MAR train & MNAR test settings. This indicates that the missingness mechanism in the train set can significantly affect model fairness on the MNAR test set. Additionally, miss-forest and auto-ml proved to be the most robust MVI techniques against test set missingness shifts under all conditions.
@@
Figure 48. True Negative Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
@@
Figure 49. Disparate Impact of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
@@
Figure 50. Selection Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
@@
Figure 51. True Negative Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
@@
Figure 52. Disparate Impact of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
@@
Figure 53. Selection Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Appendix E Spearman Correlation by Train and Test Missingness
In Figure 11 in Section 6, we reported correlations between MVI technique, model type, test missingness and performance metrics (F1, fairness and stability), based on train missingness. In this section, we supplement this analysis by reporting correlations between MVI technique, model type and performance metrics for each combination of train and test missingness.
Figure 54 is for MCAR train and different test missingness, Figure 55 is for MAR train and different test missingness, and Figure 56 is for MNAR train and different test missingness.
(a) MCAR train, MCAR test
(b) MCAR train, MAR test
(c) MCAR train, MNAR test
Figure 54. Spearman correlation ($\rho$) between MVI technique, model type, test missingness and performance metrics (F1, fairness and stability) under MCAR train and different test missingnesses. TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$
(a) MAR train, MCAR test
(b) MAR train, MAR test
(c) MAR train, MNAR test
Figure 55. Spearman correlation ($\rho$) between MVI technique, model type, test missingness and performance metrics (F1, fairness and stability) under MAR train and different test missingnesses. TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$
(a) MNAR train, MCAR test
(b) MNAR train, MAR test
(c) MNAR train, MNAR test
Figure 56. Spearman correlation ($\rho$) between MVI technique, model type, test missingness and performance metrics (F1, fairness and stability) under MNAR train and different test missingnesses. TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$